OpenStudy (anonymous):

solve 5x^2 + 3x - 4=0 by completing the square

OpenStudy (anonymous):

no

OpenStudy (anonymous):

where are you getting c and b from?

OpenStudy (anonymous):

is it -3/10

OpenStudy (anonymous):

will you stop calling me bro?

OpenStudy (anonymous):

Your original equation is set up in order of the formula he gave you. That is how you find c and b.

OpenStudy (anonymous):

http://www.purplemath.com/modules/sqrquad.htm this is a step-by-step guide. Hope it helps.

OpenStudy (anonymous):

well im not sure what to do then because thats how my online teacher said how to work it

OpenStudy (anonymous):

@stevench That website should help, it gives step-by-step. You will go to the guide near the bottom of the page, not the first one. it is pretty visual so you know what comes from where. :)

OpenStudy (anonymous):

However, I'm not so sure you can do it.. because through the steps if you divide 3 by two and square it it comes out to 2.25. Is it a direct method your teacher said or is it what the assignment requested? I don't think completing the square is the best method.. I really think factoring would work much better.

OpenStudy (anonymous):

we want to complete the square so we are going to work with \[5x ^{2}+3x\] 3 = 2(1/2) and \[\frac{ 1 }{ 2 }\div \sqrt{5}=\frac{ \sqrt{5} }{ 10}\] so 5x^{2}+3x=4\] \[5x ^{2}+3x-4=0 \rightarrow (x \sqrt{5)^{2}}+2(x \sqrt{5})(\frac{ 3\sqrt{5} }{ 10})+(\frac{ 3\sqrt{5} }{ 10})^{2}=4+(\frac{ 3\sqrt{5} }{ 10})^{2}\] then we have \[5x ^{2}+3x-4=0 \rightarrow (x \sqrt{5}+\frac{ 3\sqrt{5} }{ 10 })^{2}=4+(\frac{ 3\sqrt{5}}{ 10 })^{2}\] with this you can solve the equation easily as we've completed the squares. I think It's pretty messy to use the completion of square method on this kind o equation. I hope o helped you out. Leave a good feedback. Thanks